# Illicit Promotion on Twitter

### v2

@@ Abstract.
In this paper, we present an extensive study of the promotion of illicit goods and services on Twitter, a popular online social network(OSN). This study is made possible through the design and implementation of multiple novel tools for detecting and analyzing illicit promotion activities as well as their underlying campaigns. As the results, we observe that illicit promotion is prevalent on Twitter, along with noticeable existence on other three popular OSNs including Youtube, Facebook, and TikTok. Particularly, 12 million distinct posts of illicit promotion (PIPs) have been observed on the Twitter platform, which are widely distributed in 5 major natural languages and 10 categories of illicit goods and services, e.g., drugs, data leakage, gambling, and weapon sales. What are also observed are 580K Twitter accounts publishing PIPs as well as 37K distinct instant messaging (IM) accounts that are embedded in PIPs and serve as next hops of communication, which strongly indicates that the campaigns underpinning PIPs are also of a large scale. Also, an arms race between Twitter and illicit promotion operators is also observed. On one hand, Twitter is observed to conduct content moderation in a continuous manner and almost 80% PIPs will get gradually unpublished within six months since posted. However, in the meantime, miscreants adopt various evasion tactics to masquerade their PIPs, which renders more than 90% PIPs keeping hidden from the detection radar for two months or longer.
* These authors contributed equally to this research.
copyright: noneconference: ACM Conference on Computer and Communications Security; Due 14 Jan 2022; Los Angeles, U.S.Ajournalyear: 2022
@@ 1. Introduction
Traditionally, illicit goods and services are promoted either offline or through anonymous online marketplaces (Soska and Christin, 2015), e.g., the Silk Road that runs as an onion service. However, these channels tend to have a very constrained audience base (Christin, 2013), and are thus not feasible to promote illicit products to large-scaled regular online users. To reach a wider customer base especially regular online users, alternative promotion techniques have been developed and adopted by miscreants. One typical example is the search engine poisoning attack (John et al., 2011; Liao et al., 2016a) which involves the injection of illicit promotion into benign but vulnerable websites as well as misleading search engines to index the poisoned webpages with high page rank. However, since a compromised website can be recovered quickly (Leontiadis et al., 2014), the miscreants have to continuously identify and compromise new websites so as to maintain the magnitude of their promotion campaigns.
@@
Figure 1. Exemplary posts of illicit promotions on Twitter.
@@
Instead, illicit promotion on online social networks (OSNs) is traditionally considered as either infeasible or uncommon, since OSNs typically enforce strict content moderation against accounts and posts. However, it is not the case anymore. In this study, we observe that posts of illicit promotion (PIPs) are being distributed at a concerning scale on Twitter, a major online social network platform. Figure 1 presents two PIP examples published on the Twitter platform. One post (Figure1(a)) is intended to promote drug trading. In the attached image, we can see some hand-written Thai words and several bags of products that look like heroin. And the other post (Figure1(b)) advertises the service of forging certificates and photo IDs. In this post, the main text is masqueraded as a benign English sentence, while both the Chinese username and the attached images clearly promotes a fake certificate service. Motivated by such examples, this study aims to gain an in-depth understanding of illicit promotion on Twitter. And, we focus on the following research questions. First of all, what illicit goods and services are being promoted on Twitter? Besides, how can posts of illicit promotion evade content moderation of Twitter and get published? Also, how do the underlying operators of illicit goods and services communicate with the victims (potential customers) as exposed to their PIPs?
@@
To fulfill these research questions, multiple technical challenges must be tackled. Firstly, given a limited access to the Twitter platform, it is challenging to identify PIPs with a good coverage. Also, there are no existing tools that can accurately distinguish PIPs from benign Twitter posts (i.e., tweets), not to mention classifying PIPs into well-known categories of illicit goods and services. Furthermore, cybercrime operators tend to embed in PIPs the contacts to facilitate the next-step communication with victims. However, such contacts are of diverse categories and are typically presented in an evasive manner that renders automatic recognition error-prone while keeping them still human-readable. We have addressed these technical challenges through two novel tools. One is the PIP hunter which can not only efficiently search the Twitter platform with known PIP keywords (e.g., relevant hashtags), but also classify whether a given tweet is a PIP or not using machine learning, as well as snowballing this hunting process by automatically generating novel PIP-relevant keywords. The other tool is designed to gain an in-depth understanding of captured PIPs and their underlying campaigns. We thus name it as the PIP analyzer. It consists of a multiclass classifier to classify PIPs into well-defined categories of illicit goods and services, an PIP contact extractor based on named entity recognition (NER), as well as a clustering module designed to group PIPs into their underlying campaigns of illicit goods and services. Leveraging this novel toolchain, we have conducted an extensive study on illicit promotion on Twitter. Below, we highlight the key findings of this study.
@@
First of all, illicit promotion on Twitter features a large scale, diverse categories and products, and a wide distribution across natural languages and Twitter accounts. Specifically, in total, we have captured 12,401,082 distinct PIPs as well as 580,530 Twitter accounts. These Twitter accounts either publish PIPs (i.e., PIP accounts) or promote illicit goods and services in their personal profiles. Besides, The captured PIPs are widely distributed in multiple natural languages and reside in 10 well-defined categories of illicit goods and services. The top categories with most PIPs are porn & sex services (69.78%), gambling (13.34%), illegal drugs (8.04%), money-laundry (4.09%) and data leakage(2.03%). Also, from PIPs in each category, a diverse set of specific products have been observed, e.g., methamphetamine and marijuana in the category of illegal drug, and ID cards and passports in the category of data theft and leakage.
Besides, illicit promotion is not limited to Twitter, but turns out to be a cross-OSN issue. Particularly, we have observed the extensive distribution of PIPs on other three popular OSN platforms including Facebook, TikTok, and Youtube. Particularly, when searching these OSNs with sampled PIP keywords that cover all the ten categories, 35% keywords have one or more PIPs returned as top search results on Youtube, while it is 83% on Facebook, and 52% on TikTok. Also, some PIPs observed across OSNs have the same contacts embedded, which suggests that they belong to the same underlying campaign and some miscreants are promoting their illicit goods and services across OSN platforms. Such an observation also highlights the necessity of cross-OSN collaboration in terms of mitigating illicit promotion.
@@
Then, an arms race is observed between illicit promotion campaigns and Twitter content moderation. On one hand, various evasion techniques have been adopted by PIP operators, e.g., the use of various jargon words, composing PIPs with multilingual characters, and masquerading the tweet text as benign but injecting illicit promotion elements into usernames, media files, or even poll options. The adoption of diverse evasion techniques may explain why over 90% PIPs can survive the first two months since published. On the other hand, Twitter carries out continuous content moderation and almost 80% PIPs have got banned six months later after being published, as learned through periodically revisiting captured PIPs and checking their availability.
@@
Furthermore, when it comes to further communication with victims exposed to PIPs, most PIP operators prefer instant messaging platforms rather than Twitter itself, especially end-to-end encrypted ones. As the results, we have extracted from PIPs, 37,621 instant messaging accounts which include 9,644 Telegram accounts, 11,561 WeChat accounts, 12,702 QQ accounts, 225 WhatsApp accounts, and 3,489 LINE accounts. Also, miscreants underpinning PIPs of different categories appear to vary a lot in terms of their preferred instant messaging platform, e.g., operators of money laundering or weapon sales prefer Telegram accounts while data leakage services prefer WeChat.
Our contributions can be summarized as below.
$\bullet$ To the best of our knowledge, this is the first extensive security study of illicit promotion on the Twitter platform, which have distilled a set of previously unknown security findings.
@@
$\bullet$ Two novel tools have been designed and implemented to capture and analyze illicit promotion, along with a large dataset of PIPs and PIP contacts captured. We have made the tools and datasets available to the public to promote research in this domain [^1] Available in https://illicit-promotion.github.io/.
@@ 2. Background and Related Works
The promotion and communication of the underground economy. The underground economy, or the shadow economy refers to the production, promotion, trade, and distribution of goods or services that are deemed illicit or even illegal. It ranges from traditional categories (such as drug trading and child pornography), to new ones such as hacking services and the trading of illegal data (Du et al., 2018). In many studies, the underground economy is called cybercrime since it has many activities that either have moved to the Internet or owe their existence to the Internet, and we use both terms interchangeably in this study.
@@
A long line of studies have profiled the underground economy from various aspects, among which, a large portion are dedicated to a single cybercrime category, e.g., counterfeit or unlicensed pharmaceuticals (Kanich et al., 2011; Leontiadis et al., 2011; McCoy et al., 2012), drug trading (Aldridge and Askew, 2017; Li et al., 2021), illegal online gambling (Yang et al., 2019; Gao et al., 2021), malware distribution (Caballero et al., 2011; Kotzias et al., 2016; Thomas et al., 2016; Kotzias et al., 2021), among others. Another line of work examines holistic infrastructures that promote diverse categories of illicit goods and services. Through these studies, various promotion channels have been identified and profiled. One prominent example is the anonymous markets that run as Tor hidden services. Christin et al. (Christin, 2013) conducted a measurement study between 2011 and 2012 on such an anonymous marketplace, namely the Silk Road. The Silk Road marketplace was found to be mostly used in the trading of controlled substances and narcotics. Also, most items were on sale for less than three weeks, and only 564 sellers and 24,000 unique items were observed. Furthermore, Soska et al. (Soska and Christin, 2015) moved forward and profiled the longitudinal evolution of anonymous marketplaces between 2013 and 2015. Still, most sales were found to be drug-related.
To reach a wider customer base, especially regular Internet users, miscreants have also abused or even compromised popular online services. One typical and well-studied example is promotional infections or search poisoning attacks (John et al., 2011; Leontiadis et al., 2014; Liao et al., 2016a) wherein the attacker compromises a legitimate website, injects promotional and harmful webpages, induce the search engines to index these webpages with high page ranks, and ultimately expose benign search users to the injected harmful webpages. Such promotional infections have been used to promote diverse cybercrime activities, e.g., online gambling (Liao et al., 2016a), unlicensed pharmacies (Leontiadis et al., 2014), etc. As revealed by (Leontiadis et al., 2014), the median time to recover such promotional infections is around 15 days. To detect promotional infections, John et al. (John et al., 2011) proposed the detection of infected webpages by looking at its URL parameters instead of visiting the webpage. Besides, Liao et al. (Liao et al., 2016a) utilized the semantic inconsistency between the injected cybercrime content and the legitimate context of the infected website to decide if a webpage is a promotional infection or not.
@@
Besides, to gain a deep understanding of the underground economy, an important obstacle exists in the use of jargon words among sellers and buyers of illicit goods and services. Some research efforts (Zhao et al., 2016; Yang et al., 2017a; Yuan et al., 2018; Zhu et al., 2021) have thus been invested to explore how to identify and understand such kinds of jargon words. Particularly, Yang et al (Yang et al., 2017a) explored identifying jargon words from search engine keywords as promoted in black hat SEO campaigns, while Yuan et al. (Yuan et al., 2018) detected whether a word is a jargon or not by comparing the semantic discrepancy between cybercrime contexts and benign ones for a given word.
Moving forward from these studies on the the underground economy, we have extensively explored, for the first time, how miscreants promote their illicit goods and services on popular online social networks, as well as how to address this concerning online abuse problem.
Spam in online social networks. The widespread adoption of online social networks (OSN) is paralleled by various spamming activities which include but not limited to unsolicited advertisements, scams, malware distribution spams, etc. To fight against such spamming activities, different detection methods have been developed and evaluated to detect spam messages (Gao et al., 2012; Xu et al., 2016; Zheng et al., 2019), spam accounts (spammers) (Benevenuto et al., 2010; Yang et al., 2013; Miller et al., 2014), and spam URLs (Thomas et al., 2011a). Particularly, Gao et al. (Gao et al., 2012) considered spam detection as a two-step task in which messages are first grouped into clusters and the resulting message clusters are further classified as spam campaigns or not. Instead, Miller et al. (Miller et al., 2014) adopted anomaly detection methodologies for spam detection. In addition to spam detection, many studies (Grier et al., 2010; Thomas et al., 2011b, 2013) have also distilled novel findings regarding spam activities on OSNs. Particularly, Grier et al. (Grier et al., 2010) revealed in 2010 the distribution of spam URLs on Twitter, and found that 8% URLs posted on Twitter pointed to malicious websites (e.g., malware, and scams) which however had a clickthrough rate as low as 0.13% due to the anti-spam measures of Twitter. Besides, Thomas et al. (Thomas et al., 2013) uncovered the trade of fraudulent Twitter accounts in the underground economy and the important role that the accounts had played in spamming activities on Twitter.
Different from these studies, this study focuses on illicit promotions on OSNs (especially Twitter). Also, In contrast to OSN spamming, illicit promotion aims to advertise illicit goods/services rather than phishing or scamming the victims.
@@
Illicit promotion on OSNs. Previous works have also made efforts to reveal illicit promotion on OSNs, but most of them focus on a specific category or entity instead of comprehensive detection and understanding. For example, authors of (Kalyanam and Mackey, 2017; Katsuki et al., 2015) paid attention to illicit pharmacies on Twitter and (Yuan et al., 2019) studied adversarial pornography images crawled from popular OSNs along with the underground business behind them. In this paper, we step forward to build up a general tool chain to detect and analyze PIPs along with efficacy demonstrated for diverse PIP categories and multilingual PIP instances. Furthermore, most findings distilled from this study are applicable to the ecosystem of PIPs rather than being limited to a specific PIP category.
Language models. A language model (langmodel, 2023) is a probability distribution over a set of natural language words, and many NLP tasks use DNN-based language models like RNN-based word2vec (Mikolov et al., 2013) and transformer-based BERT (Devlin et al., 2018). Such a DNN-based language model is usually trained on a large corpus of unlabeled text documents through self-supervised training tasks such as masked language modeling, i.e., predicting a missing word in a given sentence. Since the hidden layers of such language models can well represent natural language words in a fixed-size and high-dimensional vector space and are thus commonly used for text encoding, i.e., word embedding. However, training a large and general-purpose language model is both time-consuming and computing-intensive, which motivates the emergence and increasing adoption of the paradigm of pre-training and fine-tuning. In this paradigm, to fulfill a NLP task, rather than building up a DNN model from scratch, a general-purpose and pre-trained language model is first adopted and then fine-tuned on a labeled dataset that is specific to the given NLP task and can be small-scaled. This paradigm outperforms numerous NLP tasks (Devlin et al., 2018; Koroteev, 2021; González-Carvajal and Garrido-Merchán, 2020).
In this study, to facilitate the detection and understanding of PIPs, we have adopted the paradigm of pre-training and fine-tuning in multiple text classification tasks, e.g., binary PIP classification, multiclass PIP classification, and PIP contact recognition, as detailed in §3. Particularly, across these NLP tasks, we adopt the multilingual BERT (mul, 2023) as the pre-trained language model, which is built up through two self-supervised training tasks, namely, masked language modeling and next sentence prediction, upon unlabeled Wikipedia corpus which contain numerous text documents written in 102 natural languages.
@@ 3. Methodology
In this section, we present the details of our methodology to capture and analyze posts of illicit promotion (PIPs) on Twitter. As illustrated in Figure 2, this methodology is comprised of two key modules. One is the PIP hunter, an automatic pipeline to capture illicit promotion tweets as well as collecting the relevant Twitter accounts, as detailed in §3.1. The other is the PIP analyzer3.2), which is designed to profile PIPs with regards to their categories of illicit services and goods, next-hop contacts, and the underlying campaigns.
Figure 2. The methodology to capture and analyze posts of illicit promotion (PIPs) on Twitter.

### v3

# Detecting and Understanding the Promotion of Illicit Goods and Services on Twitter

@@ Abstract.
In this study, we reveal, for the first time, popular online social networks (especially Twitter) are being extensively abused by miscreants to promote illicit goods and services of diverse categories. This study is made possible by multiple machine learning tools that are designed to detect and analyze Posts of Illicit Promotion (PIPs) as well as revealing their underlying promotion campaigns. Particularly, we observe that PIPs are prevalent on Twitter, along with extensive visibility on other three popular OSNs including YouTube, Facebook, and TikTok. For instance, applying our PIP hunter to the Twitter platform for 6 months has led to the discovery of 12 million distinct PIPs which are widely distributed in 5 major natural languages and 10 illicit categories, e.g., drugs, data leakage, gambling, and weapon sales. Along the discovery of PIPs are 580K Twitter accounts publishing PIPs as well as 37K distinct instant messaging accounts that are embedded in PIPs and serve as next hops of communication with prospective customers. Also, an arms race between Twitter and illicit promotion operators is also observed. Especially, 90% PIPs can survice the first two months since getting published on Twitter, which is likely due to the diverse evasion tactics adopted by miscreants to masquerade PIPs.
Illicit Promotion; Online Abuse; Cybercrime; Twitter; Online Social Networks
copyright: acmlicensedjournalyear: 2025doi: 10.1145/10.1145/3696410.3714550booktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28–May 2, 2025, Sydney, NSW, Australiaisbn: 979-8-4007-1274-6/25/04conference: Proceedings of the ACM Web Conference 2025; April 28–May 2, 2025; Sydney, NSW, Australia.ccs: Security and privacy Social network security and privacy
@@ 1. Introduction
Traditionally, illicit goods and services are promoted either offline or through anonymous marketplaces (Soska and Christin, 2015), e.g., the Silk Road running as an onion service. However, these channels tend to have a very constrained audience base (Christin, 2013), and are thus not feasible to promote illicit products to large-scale regular online users. To reach a wider customer base, especially regular online users, alternative promotion techniques have been developed and adopted by miscreants. One typical example is the search engine poisoning attack (John et al., 2011; Liao et al., 2016a) which involves the injection of illicit promotion into benign but vulnerable websites as well as misleading search engines to index the poisoned webpages with high page rank. However, since a compromised website can be recovered quickly (Leontiadis et al., 2014), the miscreants have to continuously identify and compromise new websites so as to maintain the magnitude of their promotion campaigns.
@@
Figure 1. Example posts of illicit promotion on Twitter.
@@
Instead, illicit promotion on online social networks (OSNs) is traditionally considered either infeasible or uncommon, since OSNs typically enforce strict content moderation against accounts and posts. However, this is not the case anymore. In this study, we observe that posts of illicit promotion (PIPs) are being distributed at a concerning scale on Twitter, a major online social network platform. Figure 1 presents two PIP examples published on the Twitter platform. One post (Figure1(a)) is intended to promote drug trading, whose attached image shows some hand-written Thai words and several bags of products that look like heroin. And the other post (Figure1(b)) advertises the service of forging certificates and photo IDs. In this post, the main text is masqueraded as a benign English sentence, while both the Chinese username and the attached images clearly promote a fake certificate service. Motivated by such examples, this study aims to gain an in-depth understanding of illicit promotion on Twitter. We focus on the following research questions. First of all, what illicit goods and services are being promoted on Twitter? Besides, how can posts of illicit promotion evade content moderation of Twitter and get published? Also, how do the underlying operators of illicit goods and services communicate with the victims (potential customers) as exposed to their PIPs?
@@
To fulfill these research questions, multiple technical challenges must be tackled. Firstly, given limited access to the Twitter platform, it is challenging to identify PIPs with good coverage. Also, there are no existing tools that can accurately distinguish PIPs from benign Twitter posts (i.e., tweets), not to mention classifying PIPs into well-known categories of illicit goods and services. Furthermore, cybercrime operators tend to embed in PIPs the contacts to facilitate the next-step communication with victims. However, such contacts are of diverse categories and are typically presented in an evasive manner that renders automatic recognition error-prone while keeping them still human-readable. We have addressed these technical challenges through two novel tools. One is the PIP hunter which can not only efficiently search the Twitter platform with known PIP keywords (e.g., relevant hashtags), but also classify whether a given tweet is a PIP or not using machine learning, as well as snowballing this hunting process by automatically generating novel PIP-relevant keywords. The other tool is designed to gain an in-depth understanding of captured PIPs and their underlying campaigns. We thus name it as the PIP analyzer. It consists of a multiclass classifier to classify PIPs into well-defined categories of illicit goods and services, a PIP contact extractor based on named entity recognition (NER), as well as a clustering module designed to group PIPs into their underlying campaigns of illicit goods and services. Leveraging this novel toolchain, we have conducted an extensive study on illicit promotion on Twitter. Below, we highlight the key findings of this study.
@@
Findings. First, illicit promotion on Twitter features a large scale, diverse categories and products, and a wide distribution across natural languages and Twitter accounts. Specifically, in total, we have captured 12,401,082 distinct PIPs as well as 580,530 Twitter accounts. These Twitter accounts either publish PIPs (i.e., PIP accounts) or promote illicit goods and services in their personal profiles. Besides, the captured PIPs are widely distributed in multiple natural languages and reside in 10 well-defined categories of illicit goods and services. The top categories with the most PIPs are porn & sex services (69.78%), gambling (13.34%), illegal drugs (8.04%), money-laundry (4.09%) and data leakage(2.03%). Also, from PIPs in each category, a diverse set of specific products have been observed, e.g., methamphetamine and marijuana in the category of illegal drug, and ID cards and passports in the category of data theft and leakage.
@@
Then, an arms race is observed between illicit promotion campaigns and Twitter content moderation. On one hand, various evasion techniques have been adopted by PIP operators, e.g., the use of various jargon words, composing PIPs with multilingual characters, and masquerading the tweet text as benign but injecting illicit promotion elements into usernames, media files, or even poll options. The adoption of diverse evasion techniques may explain why over 90% of PIPs can survive the first two months since published. On the other hand, Twitter carries out continuous content moderation and almost 80% of PIPs have got banned six months later after being published, as learned through periodically revisiting captured PIPs and checking their availability.
@@
Furthermore, for further communication with prospective customers, most PIP operators prefer instant messaging platforms rather than Twitter itself, especially end-to-end encrypted ones. Especially, we have extracted from PIPs over 37K accounts of five instant messaging platforms including Telegram, WeChat, QQ, WhatsApp, and LINE. Also, miscreants underpinning PIPs of different categories appear to vary a lot in terms of their preferred instant messaging platform, e.g., operators of money laundering or weapon sales prefer Telegram accounts while data leakage services prefer WeChat.
@@
Contributions. Our key contributions are two-fold. On one side, to the best of our knowledge, this is the first extensive security study of illicit promotion on the Twitter platform, which has distilled a set of previously unknown security findings. On the other side, two novel tools have been designed and implemented to capture and analyze illicit promotion, along with a large dataset of PIPs and PIP contacts captured. Both these tools and datasets have been made available at https://illicit-promotion.github.io.
@@ 2. Background and Related Works
The underground economy. The underground economy, or the shadow economy refers to the production, promotion, trade, and distribution of goods or services that are deemed illicit or even illegal. It ranges from traditional categories (e.g., drug trading and child pornography) to new ones such as hacking services and the trading of illegal data (Du et al., 2018). In many studies, the underground economy is called cybercrime since it has many activities that either have moved to the Internet or owe their existence to the Internet, and we use both terms interchangeably in this study.
@@
Among studies profiling the underground economy, a large portion is dedicated to a specific cybercrime category, e.g., counterfeit or unlicensed pharmaceuticals (Kanich et al., 2011; Leontiadis et al., 2011; McCoy et al., 2012), drug trading (Aldridge and Askew, 2017; Li et al., 2021), illegal online gambling (Yang et al., 2019; Gao et al., 2021), malware distribution (Caballero et al., 2011; Kotzias et al., 2016; Thomas et al., 2016; Kotzias et al., 2021), among others. Another line of work examines holistic infrastructures that promote diverse categories of illicit goods and services. Through these studies, various promotion channels have been identified and profiled. To reach a wider customer base, especially regular Internet users, miscreants have also abused or even compromised popular online services. A well-studied example is search engine poisoning attacks (John et al., 2011; Leontiadis et al., 2014; Liao et al., 2016a) (see Appendix A).
@@
Besides, to gain a deep understanding of the underground economy, an important obstacle exists in the use of jargon words among sellers and buyers of illicit goods and services. Some research efforts (Zhao et al., 2016; Yang et al., 2017a; Yuan et al., 2018; Zhu et al., 2021) have thus focused on identifying and understanding such kinds of jargon words. Particularly, Yang et al (Yang et al., 2017a) explored identifying jargon words from search engine keywords as promoted in black hat SEO campaigns, while Yuan et al. (Yuan et al., 2018) detected jargon by analyzing the semantic discrepancy between cybercrime contexts and benign ones for a given word.
@@
Illicit promotion on OSNs. Previous works have also made efforts to reveal illicit promotion on OSNs, but most of them focus on a specific category or entity instead of comprehensive detection and understanding. For example, authors of (Kalyanam and Mackey, 2017; Katsuki et al., 2015) paid attention to illicit pharmacies on Twitter and (Yuan et al., 2019) studied adversarial pornography images crawled from popular OSNs along with the underground business behind them. In this paper, we step forward to build up a general tool chain to detect and analyze PIPs along with efficacy demonstrated for diverse PIP categories and multilingual PIP instances. Furthermore, most findings distilled from our study are applicable to the ecosystem of PIPs rather than being limited to a specific PIP category.
@@ 3. Methodology
Figure 2. The methodology to capture and analyze PIPs.
In this section, we present the methodology to capture and analyze posts of illicit promotion (PIPs) on Twitter. As illustrated in Figure 2, this methodology is comprised of two key modules. One is the PIP hunter3.1), an automatic pipeline to capture posts of illicit promotion and relevant Twitter accounts. The other is the PIP analyzer3.2), which is designed to profile PIPs with regards to categories, next-hop contacts, and the underlying campaigns.